OpenStudy (anonymous):

Each child ticket for a ride costs $2, while each adult ticket costs $6. If the ride collected a total of $148, and 38 tickets were sold, how many of each type of ticket were sold? 20 child and 18 adult 19 child and 19 adult 14 child and 20 adult 148 adult and 38 child

OpenStudy (texaschic101):

c + a = 38 (child tickets + adult tickets = 38 , because there was 38 tickets sold) 2c + 6a = 148 c = 38 - a now sub 38 - a in for c in the other equation 2c + 6a = 148 2(38 - a) + 6a = 148 76 - 2a + 6a = 148 -2a + 6a = 148 - 76 4a = 72 a = 18 now sub 18 in for a in either of the original equations c + a = 38 c + 18 = 38 c = 38 - 18 c = 20 check... 2c + 6a = 148 2(20) + 6(18) = 148 40 + 108 = 148 148 = 148 (correct) sold 18 adult tickets and 20 child tickets any questions ?

OpenStudy (anonymous):

None at all, thank you for explaining how you got to the answer! i really Appreciate it!

OpenStudy (texaschic101):

no problem :)